OpenStudy (anonymous):

Pre Calc: Write an equation in slope-intercept form for the line passing through A(–2, 1) and having a slope of 0.5.

OpenStudy (anonymous):

slope intercept form is y=mx+b

OpenStudy (anonymous):

your slope (1/2) is m. so y = 1/2x + b now plug in your x and y and solve for b

OpenStudy (anonymous):

so it would be 1 = (1/2)(-2) + b

OpenStudy (anonymous):

Ya that is one way u can do it...... @samalami Or u can use direct point-slope formula.... that u know

OpenStudy (anonymous):

Thank you so much!! @ksandoval @sauravshakya

OpenStudy (anonymous):

sure! (:

OpenStudy (anonymous):

welcome